40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  TCS elog  Not logged in ELOG logo
Message ID: 254     Entry time: Fri Jul 8 07:06:41 2022
Author: Jc 
Type: Misc 
Category: VAC 
Subject: Dewar 

[Stephen, JC]

Stephen and I added a new valve to Dewar's vacuum system. This valve allows the flow of atmospheric pressure into the system. We added 3 components to the system which were the valve, an adapter, and a T-intersection. After added these components, we continued to pump down only the highlighted yellow and green area with Dewar being closed off. The system pumped down to .1 mTorr until we decided to close off the pump. Once we close the pump off, we noticed the pressure began to rise. We took apart the system again and looked over the O-rings. We came across one ring with a sticky and clear material surround the rim and another ring with a fiber on it. We proceeded to clean and reassemble the system, but ran into this same issue. We tried to find where this leak was by squirting isopropyl around each ring and posibbly so a slower rise in pressure, but had no luck.

Afterwards, I checked both T-intersections individually and came across the same rise in pressure after closing off the valve for both tests. We suspect this may have been normal beforehand?

We continued by putting it all back together and taking data of the pumpdown.

Attachment 1: 29FCB85E-0AC8-43ED-8A43-31DC04459CD4_1_201_a.jpeg  665 kB  Uploaded Fri Jul 8 10:12:42 2022  | Hide | Hide all
29FCB85E-0AC8-43ED-8A43-31DC04459CD4_1_201_a.jpeg
ELOG V3.1.3-